UT 20130521 Technical Observing - SX AdSec Realignment

Observer: DMiller, JHill (Tucson) ASkemer, VBailey (LBTO)
Telescope Operator: DHuerta (LBTO)
AO Support: JCGuerra (Home)
Telescope Support: (LBTO)
Instrument Support: (LBTO)

Summary

LFBG

Performed the full manual alignment on LFBG
  • Preset in ACQUIRE mode, star a lower edge for acquisition image
  • Move star to center via Coma Free Pointing on Control GUI (Tip = +18", TILT = 0)
  • Remove coma in out of focus image by putting Z7 on the secondary hexapod
  • Record value (see below)

Collected field data at LFBG
  • Analyze data and find binodal
  • Adjust Coma Free Pointing to TIP= -7, TILT = +25 (delta -25, 25)

Collect second set of field data
  • binodal measured to be 0.0+-5"
  • Good.

Set Sextractor scale factor to the values measured in April.

Set WFS_rotAngleOffs to 6.0 from 9.0 from zernike crosstalk measurements

LUCI AGw#1 is released for science operation

Set zernike factors for 12-21 to 0 so ready for science

LDG

Put original SX AdSec pointing and collimation file back in use for MODS. Acquisition looks round, well collimated.

Put new LFBG pointing and collimation models into LDG. MODS well pointing and collimated

Set Sextractor scale factor to the values measured in April.

Set sextractor_wfsc.cfg BACK_SIZE to 10

Collected a field data set. 400-600 nm of correction measured off-axis. Took some MODS Blue images and see structure in images in center of chip. Not sure why. Need to analyze these field data sets to understand if there is binodal. May need to start with MODS tonight if binodal is found. TBD.

Set zernike factors for 12-21 to 0 so ready for science

Hand over to LBTI at 2:15 am UTC

LBTI@LCBG

LBTI testing (Bailey, Skemer on site, Hinz remote, JC AO support):

Pointing for LBTI station was nominal. We started with 10, 66, 153 mode reconstructors. All were stable. Seeing was 1-1.3". Tried 400 mode reconstructor. Several RIP's were caused by the same actuator. Maybe made worse by the mediocre seeing. Requested that JC remove one actuator. Successfully closed loop with 400 modes. Performance appears similar to Feb. and April observations.

Offsetting and further stability tests still needed, but the system appears usable for observations. smile

Details

AGW1@LFBG

03:15 Initial Acquire preset misses star by only ~8 arcsec at EL=76, see some coma. Will do full alignment procedure

03:20 Full Alignment
  • ptreset to put telescope back on pointing model
  • Coma free pointing to M2 only on Control GUI. Tip:+18, Tilt: 0
  • Put -700nm Z7 on M2
  • Global, -1.89, 0, 0, 0, -318.5
  • Manual, 0, -0.919, 0, 176.759, 0
  • Active, 0, 0.652, 0, -67.961, 0
  • New Global -1.89, -0.267, 0, 108.798, -318.5

03:40 preset BS9148 ACQUIRE, set IE, CA, ACTIVE

04:00 Modify LUCI_L_RM11.cfg
  • cp LUCI_L_RM_sub11.cfg LUCI_L_RM.cfg.20130326
  • Edit LUCI_L_RM_sub11.cfg: Set sextScale values to new ones measured in 201303
  • cp LUCI_L_RM_sub11.cfg LUCI_L_RM.cfg.20130521

04:05 Modify LUCI_L.cfg
  • cp LUCI_L.cfg LUCI_L.cfg.20130328
  • Edit LUCI_L.cfg: Set WFS_rotAngleOffs from 9.0 to 6.0
  • cp LUCI_L.cfg LUCI_L.cfg.20130521

04:15 Seeing jump from 1.1" to 1.5-1.9" on wfsc at wfscimage 49. Note, 4-22 all being sent

04:30 iif_preset, BS9152

04:35 field_collect, bs9152_field.list
  • data in /home/dlmiller/data/20130521/lfbg_field_1
  • seeing ~1.1
  • 04:50 10 minutes of bad seeing (2-2.5"). field_collect routine stop with failure to fit_outer_pupil, not enough SH spots. Need to put in check
  • restart field_collect at 05:02 with START=3. seeing ~1.1 again
  • restart field_collect at 05:12 with START=5. Failed again. David stopped and started GCS
  • restart field_collect at 05:15 with START=5.
  • want a TIP delta = -25
  • want a TILT delta = +25

05:54 Make binodal change
  • Send TIP= -7, TILT=0
  • Set IE from -97 to -72
  • Send TIP = -7, TILT = +25
  • Set CA from -50 to -75
  • global -1.89, 0, 0, 0, -318.5
  • manual -1.27, 0.357, 0, -68.74, -245.5
  • active -0.050, -0.255, -0.015, 26.575, -5.23
  • new global -3.20, 0.102, 0, -42, -596.2

06:00 Collect another field data set
  • data in /home/dlmiller/data/20130521/lfbg_field_2
  • seeing ~0.8
  • Tip and Tilt are both ~0.0. Should have nearly 0 binodal smile

06:40 Reconfigure for MODS

MODS1@LDG

06:50 Put pointing and collimation models back to old SX AdSec values

Thus, to revert back to the previous pt and col values:

in /PSF dir:

cp SXSMMODSCollimation.20130202.dat SXSMMODSCollimation.dat

in /PCS dir:  (**NOTE, PEPSI values have changed in PCSInstrument.conf so will edit rather than copy)

emacs -nw PCSInstrument.conf
  Change LEFTPOINTINGMODEL from SX_Gregorian.20130513.ptmod SX_Gregrorian.20130202.ptmod

07:30 Put new LUCI pointing and collimate models into MODS

in /PSF dir:

cp SXSMLUCICollimation.dat SXSMLUCICollimation.20130202.dat
emacs -nw SXSMLUCICollimation.dat
   put in global offsets:  -3.20, 0.102, 0, -42, -596.2
cp SXSMLUCICollimation.dat SXSMLUCICollimation.20130521.dat
cp SXSMLUCICollimation.20130521.dat SXSMMODSCollimation.dat

in /PCS dir:

emacs -nw PCSInstrument.conf
   change MODS and LUCI LEFTPOINTINGMODEL from SX_Gregorian.20130202.ptmod to SX_Gregorian.20130521.ptmod

cp SX_Gregorian.20130202.ptmod SX_Gregorian.20130521.ptmod
emacs -nw SX_Gregorian.20130521.ptmod
   Change IE from -97 to -72 and CA from -50 to -75

07:00 Set BACK_SIZE from 22 to 10 in /home/telescope/TCS/Configuration/GCS/20130118/left&right_sextract_wfsc.cfg. This file is used by MODS. The LUCI parameter file was already at 10.

07:50 field_collect, BS9177_new.list
  • data in /home/dlmiller/data/20130521/ldg_field_1
  • seeing ~0.8

08:25 collect mods1b.20130521.0001.fits when collection: slit mask in

08:55:35 "no guide star found in guide box...." Star found in acquisition image but offset to pickoff mirror failed

08:55 preset, BS9177, MODE=ACTIVE, GS=18, /YAXIS to collimate off-axis

09:00 mods1b.20130521.0002.fits: slit mask in

09:04 mods1b.20130521.0003.fits: y-axis = -141, Look round

09:10 mods1b.20130521.0004.fits: y = -129,x = -68: Not very round frown, sad smile

09:15 Done, David Sleep MODS, Reconfigure for LBTI

-- JohnHill - 21 May 2013
Topic revision: r6 - 22 May 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback